OpenStudy (anonymous):

please help (: 5 over 2x + 15 over 2 = 15

OpenStudy (hba):

\[\frac{ 5 }{ \frac{ 2x+15 }{ 3 } }=15,Right ?\]

OpenStudy (anonymous):

do you mean this

OpenStudy (hba):

Or is it \[\frac{ 5 }{ 2x}+\frac{ 15 }{ 2 }=15\]

OpenStudy (anonymous):